OpenStudy (anonymous):

how would you integrate sqrt(1/x^2+1/x^4)?

OpenStudy (zarkon):

\[\sqrt{\frac{1}{x^2}+\frac{1}{x^4}}=\frac{\sqrt{x^2+1}}{x^2}\] now use trig substitution